# QCPINN: Quantum-Classical Physics-Informed Neural Networks for Solving PDEs

### v5

@@ 4 Quantum-Classical Physics-Informed Neural Networks (QCPINN)
Figure 1: Architecture of the quantum-classical physics-informed neural network (QCPINN) framework. The figure illustrates the hybrid structure: a classical preprocessor encodes the inputs, which are then fed into the QNN or classical neural network (NN). The QNN is implemented as either a CV-circuit or a DV-circuit with $k$ layers. A classical postprocessor then decodes the QNN or NN outputs. The integrated loss function concurrently minimizes PDE residuals and initial/boundary condition errors, ensuring the model adheres to physical constraints throughout training. Table 1 lists the various configurations of the CV and DV-circuits studied in this work.
Table 1: This work examines different configurations of the QCPINN along with the list of PDEs used for testing and generalization.
@@
Type Options Testing Generalization
CV-circuit Measurement scheme Number, quadrature
Helmholtz,
Lid-driven cavity
Wave,
Klein-Gordon,
@@
Parameterization
Phase-free (magnitude),
Full (magnitude and phase)
(also refer to algorithm 1)
@@
Topology
Alternate, cascade, cross-mesh, layered
(also refer to Fig. 2 and table 2)

### v6

@@ 4 Quantum-Classical Physics-Informed Neural Networks (QCPINN)
Figure 1: Overview of the proposed QCPINN architecture, illustrating its hybrid structure consisting of three main components: (a) a classical preprocessor for input encoding, (b) the core processing unit using either a quantum neural network (QNN) or classical neural network (NN), and (c) a classical postprocessor for decoding the QNN or NN outputs. The QNN is implemented as either a CV or DV circuit with $k$ layers. The architecture integrates a loss function to concurrently minimize PDE residuals and enforce compliance with initial and boundary conditions, thereby ensuring adherence to physical constraints during training. The various configurations of the CV and DV circuits are detailed in Table 1.
Table 1: Various configurations of the QCPINN, along with the list of PDEs used for testing. Each configuration was evaluated against the Helmholtz and lid-driven Cavity cases, with the top-performing setups subsequently selected for further assessment on additional scenarios, including Wave, Klein-Gordon, and Convection-diffusion problems.
@@
Type Options
CV-Circuit Measurement schemes Number, Quadrature
(tested on all configurations)
Helmholtz, Lid-driven Cavity
(tested on selected configurations)
Wave, Klein-Gordon,
@@
Parameterization Phase-free (magnitude), Full (magnitude and phase) (see Algorithm 1)
@@
Topology Alternate, Cascade, Cross-mesh, Layered (see Fig. 2 and Table 2)